OpenStudy (anonymous):

sqrt{x-1}-5/\sqrt{x-1}=0

OpenStudy (anonymous):

\[\sqrt{x-1}-5/\sqrt{x-1}=0\]

OpenStudy (dangerousjesse):

Did you mean \(\huge\color{darkblue}{\sqrt{x-1}-\frac{5}{\sqrt{x-1}}=0?}\)

OpenStudy (anonymous):

yeessss !

OpenStudy (dangerousjesse):

Okay, awesome!

OpenStudy (anonymous):

i will give you a medal !

OpenStudy (dangerousjesse):

Solve for x over the real numbers: \[\large \sqrt{x-1}-\frac{5}{\sqrt{x-1}} = 0\]Isolate sqrt(x-1) to the left hand side in order to cross multiply. Add \(\large \frac{5}{\sqrt{x-1}}\) to both sides: \[\large \sqrt{x-1} = \frac{5}{\sqrt{x-1}}\]Multiply both sides by an expression to clear fractions. Cross multiply: \[\large x-1 = 5\]Solve for x. Add 1 to both sides Which leaves you with..?

OpenStudy (anonymous):

how do the \[\sqrt{x-1}\] 's cancel out ?

OpenStudy (dangerousjesse):

Because if \(\sqrt{x-1}-\frac{5}{\sqrt{x-1}}=0\), then \(\sqrt{x-1}-\sqrt{x-1}=0\) or \(\sqrt{x-1}=\frac{5}{\sqrt{x-1}}\)

OpenStudy (anonymous):

but if you cross multiply you would times